The number of times the AI algorithm is successful at detecting fake news is normally distributed with a sample mean of 900 and the sample standard deviation of 30. Assume a sample size of 100 was used.
What is the 95% confidence interval (CI) estimate for the population mean?
1. ( 894.1002 ; 905.8800)
2. (894.0474 ; 905.9526)
3. (852.8959 ; 947.1041)
4. (869.9665 ; 930.0335)
5. (895.0194 ; 904.9806)
1
Expert's answer
2020-09-27T17:55:18-0400



At the 95% confidence interval, the number of times the algorithm succeeds is between 2.5% and 97.5%


"\bar x = 900" , "\sigma=30" and "n=100"


On the standard normal distribution;



"Z\ score\ at\ 2.5\% = -1.96""Z\ score\ at\ 97.5\% = 1.96"


"Z\ score =\frac{x - \bar x}{\frac{\sigma}{\sqrt{n}}}"

Therefore:


"x =\frac{\sigma}{\sqrt{n}} * Z\ score +\bar x"

The values of "x" at the given "Z\ score \ value" form the boundaries for the confidence interval


The lower boundary:



"x_{lower}=\frac{30}{\sqrt{100}} * (- 1.96) + 900 =894.1200"

"x_{upper}=\frac{30}{\sqrt{100}} * 1.96 + 900 = 905.8800"

answer: the confidence interval is

"[894.1200,\ 905.8800]"